OpenStudy (anonymous):

Suppose X and Y are independent random variables. The variance of X is equal to 16; and the variance of Y is equal to 9. Let Z = X - Y. What is the standard deviation of Z?

OpenStudy (anonymous):

SD Z = SD X-Y = sqrt(Var X-Y) = sqrt(VarX +Y) = sqrt(Var x + Var Y) = sqrt 16+9 =5